The last dividend paid by Wilden Corporation was $3.25. The dividend growth rate is expected to be constant at 1.5% for 2 years, after which dividends are expected to grow at a rate of 8.0% forever. The firm's required return (rs) is 12.0%. What is the best estimate of the current stock price?

a. $80.34
b. $81.67
c. $79.01
d. $77.68
e. $76.35
Answer #1

D0 =3.25
D1=3.25*(1+1.5%)
D2=3.25*(1+1.5%)^2
D3 =3.25*(1+1.5%)^2*(1+8%)
Terminal value =D3/(Required -Growth) =3.25*(1+1.5%)^2*(1+8%)/(12%-8%)=90.4022
Current stock Price =3.25*(1+1.5%)/(1+12%)+3.25*(1+1.5%)^2/(1+12%)^2+90.4022/(1+12%)^2=77.68(Option d is correct option)